# Cross-Domain Few-Shot Learning with Coalescent Projections and Latent Space Reservation

### v1

@@ 4.2.2 Novel Class Generation in Input Space
5 Time Complexity Analysis
@@
The crucial parts of a ViT are the Multi-Head Self-Attention (MHSA), MLPs in the transformer blocks, and the patch embedding convolution. Convolution layer is equivalent to a linear layer and is negligible in comparison to the overall calculations in $L$ blocks. The patch embedding layer generates $A=\frac{HW}{P^{2}}+1$ tokens (1 for the CLS token), where $H$, $W$, and $P$ are the image height, width, and patch size, respectively. The QKV requires $O(Ad^{2})$ operations. An ordinary attention mechanism requires $O(dA^{2}+Ad^{2})$. With CP, we have $O(2*dA^{2}+Ad^{2})=O(dA^{2}+Ad^{2})$, which does not change the time-complexity in the Big O terms. Finally, the MLPs require $O(L_{\text{MLP}}Ad_{\text{MLP}}^{2})$. Since the MLPs are shallow in practice ($L_{\text{MLP}}=2$), it becomes $O(Ad_{\text{MLP}}^{2})$
@@
Overall, the total time-complexity over $L$ transformer blocks is $O\left(L(dA^{2}+Ad^{2}+Ad_{\text{MLP}}^{2})\right)$, which is equal to the time-complexity of the ViT.
6 Experiments
In this section, we measure the effectiveness of CPLSR in practice on the BSCD-FSL benchmark [8]. Furthermore, we perform an ablation study to demonstrate the effectiveness of each component of our method.
6.1 Datasets
We measure the effectiveness of CPLSR comprehensively on the BSCD-FSL benchmark [8], which has four target datasets: ChestX [31], ISIC [4], EuroSAT [10], and CropDisease [20] by having the Mini-ImageNet [28] or the Tiered-ImageNet [23] as the base datasets. We train our model on $64$ classes from the training subset of Mini-ImageNet and $351$ classes from the training subset of Tiered-ImageNet for $1,000$ episodes. Moreover, we utilize $16$ classes from the validation subset of the Mini-ImageNet and $97$ classes from the validation subset of the Tiered-ImageNet for validation in $600$ episodes. Finally, we tested our model on the target datasets for $1,000$ episodes, and report the average top-1 accuracy with a $95\%$ confidence interval.
6.2 Baseline Methods
We compare our method with single-source inductive methods that can be reproduced. We compared our propose method with the StyleAdv [7], Wave-SAN [6], LRP [25], ATA [30], AFA [12], FWT [26], and DINO for comparison. For the Tiered-ImageNet dataset, we compare our method with the main competitors, FAP, PMF, StyleAdv, and DINO.
6.3 Implementation details
In our experiments, following the PMF [11] and StyleAdv [7] studies, we utilize the same ViT-S/16 model, which was pre-trained on ImageNet1K with DINO. The image size is 224. The experiments are performed on a single NVIDIA GeForce RTX 4090. We use the CP in all layers of the frozen ViT-S, and we utilize AdamW with a learning rate of $1\times 10^{-5}$ to optimize the CP prompt parameters. Finally, we set the $\varepsilon$ to 0.02, $E_{\text{P}}$ to $100$, and $N_{0}$ to $2$.
6.4 Results
Tab. 1 and Tab. 2 show the numerical results for all target datasets when the Mini-ImageNet and Tiered-ImageNet dataset is the base dataset, respectively. The overall results show that CPLSR outperforms all SOTA methods on both 1-shot and 5-shot settings across the two base datasets, based on the average accuracy over four datasets. In the following, we analyze the numbers in more detail.
The most challenging 1-shot results for the Mini-ImageNet experiments indicate that the proposed method outperforms the other methods on the ChestX, EuroSAT, and CropDisease datasets. In comparison to the ViT-based methods, PMF, StyleAdv, and DINO, CPLSR outperforms them on all four target datasets. In the 5-shot setting, while CPLSR beats the other methods on the EruoSAT and CropDisease, it is very close to the SOTA on ChestX.
The 1-shot results for the Tiered-ImageNet dataset show that the proposed method outperforms the SOTA methods on the ChestX, EuroSAT, and CropDisease datasets. In comparison to ViT-based methods (PMF, StyleAdv, and DINO), the proposed method is the only one that improves upon DINO’s performance. In the 5-shot setting, our method significantly outperforms the other methods on average. While CPLSR beat all methods on ChestX, EuroSAT, and CropDisease, PMF performs well on ISIC. However, PMF significantly decreases the accuracy of the model, which was pre-trained with DINO, on the other three datasets.
1-shot Arch. ChestX ISIC EuroSAT CropDisease Avg. $\downarrow$
FWT RN-10 $22.04\pm 0.46$ $31.58\pm 0.67$ $62.36\pm 1.05$ $66.36\pm 1.04$ $45.58$
AFA RN-10 $22.92\pm 0.20$ $33.21\pm 0.30$ $63.12\pm 0.50$ $67.61\pm 0.50$ $46.72$
FAP GNN $22.36\pm 0.20$ $\boldsymbol{35.63\pm 0.40}$ $62.96\pm 0.50$ $69.97\pm 0.50$ $47.73$
FAP TPN $21.56\pm 0.20$ $33.63\pm 0.40$ $62.62\pm 0.50$ $76.11\pm 0.50$ $48.48$
Wave-SAN RN-10 $22.93\pm 0.49$ $33.35\pm 0.71$ $69.64\pm 1.09$ $70.80\pm 1.06$ $49.18$
PMF + ViT-S $21.73\pm 0.30$ $30.36\pm 0.36$ $70.74\pm 0.63$ $80.79\pm 0.62$ $50.91$
LDP-Net RN-10 $22.21$ $33.44$ $73.25$ $81.26$ $52.54$
DINO ViT-S $22.92\pm 0.32$ $33.24\pm 0.44$ $73.59\pm 0.61$ $82.01\pm 0.59$ $52.94$
CPLSR ViT-S $\boldsymbol{23.00\pm 0.31}$ $33.49\pm 0.42$ $\boldsymbol{74.54\pm 0.60}$ $\boldsymbol{83.46\pm 0.58}$ $\boldsymbol{53.62}$
5-shot Arch. ChestX ISIC EuroSAT CropDisease Avg. $\downarrow$
FAP TPN $24.15\pm 0.20$ $44.58\pm 0.30$ $80.24\pm 0.30$ $88.34\pm 0.3$ $59.33$
AFA RN-10 $25.02\pm 0.20$ $46.01\pm 0.40$ $85.58\pm 0.40$ $88.06\pm 0.30$ $61.17$
FAP GNN $25.31\pm 0.20$ $47.60\pm 0.40$ $82.52\pm 0.40$ $91.79\pm 0.30$ $61.81$
LDP-Net RN-10 $26.88$ $48.44$ $84.05$ $91.89$ $62.82$
PMF ViT-S $\boldsymbol{27.27}$ $\boldsymbol{50.12}$ $85.98$ $92.96$ $64.08$
DINO ViT-S $26.86\pm 0.33$ $47.87\pm 0.46$ $89.79\pm 0.31$ $94.70\pm 0.32$ $64.81$
CPLSR ViT-S $27.14\pm 0.33$ $48.08\pm 0.47$ $\boldsymbol{90.69\pm 0.31}$ $\boldsymbol{95.11\pm 0.31}$ $\boldsymbol{65.26}$
Table 1: 5-way k-shot classification accuracy on BSCD-FSL with Mini-ImageNet as the base dataset. + shows the experiment reported by [7].
1-shot Arch. ChestX ISIC EuroSAT CropDisease Avg. $\downarrow$
FAP GNN $21.69\pm 0.23$ $\boldsymbol{34.16\pm 0.39}$ $66.11\pm 0.57$ $70.29\pm 0.56$ $48.06$
DINO ViT-S $22.92\pm 0.32$ $33.24\pm 0.44$ $73.59\pm 0.61$ $82.01\pm 0.59$ $52.94$
CPLSR ViT-S $\boldsymbol{22.97\pm 0.31}$ $33.40\pm 0.43$ $\boldsymbol{74.59\pm 0.60}$ $\boldsymbol{83.71\pm 0.58}$ $\boldsymbol{53.67}$
5-shot Arch. ChestX ISIC EuroSAT CropDisease Avg. $\downarrow$
FAP GNN $25.05\pm 0.25$ $45.16\pm 0.37$ $86.53\pm 0.35$ $90.64\pm 0.32$ $61.85$
PMF ViT-S $25.16\pm 0.43$ $\boldsymbol{49.11\pm 0.67}$ $86.29\pm 0.52$ $92.49\pm 0.48$ $63.01$
DINO ViT-S $26.86\pm 0.33$ $47.87\pm 0.46$ $89.79\pm 0.31$ $94.70\pm 0.32$ $64.81$
CPLSR ViT-S $\boldsymbol{27.01\pm 0.33}$ $48.04\pm 0.46$ $\boldsymbol{90.57\pm 0.29}$ $\boldsymbol{95.15\pm 0.30}$ $\boldsymbol{65.19}$
Table 2: 5-way k-shot classification accuracy on BSCD-FSL with Tiered-ImageNet as the base dataset. All methods are tested on our machine.
6.5 Ablation Study
Since the backbone should be frozen, the only learnable components of the network are the CPs that we used in all layers. Therefore, we consider the following cases to study: First, utilizing the CP only with pseudo-embeddings and without rotation-based augmentations, and second, utilizing the CP with rotation-based augmentations only. Finally, we consider AttnScale [1] and plain prompts with a length of two for every layer instead of using the CPs. We also display the DINO accuracies in both 1-shot and 5-shot settings to facilitate the comparison. Tab. 3 shows the results.
The results show that the CP is the crucial part of the method. By utilizing plain prompts, the network would even lose its accuracy. Moreover, while the AttnScale performs better than the plain prompts, it does not show an improvement over DINO. Regarding the augmentations, the results indicate that while utilizing the pseudo-classes has a slight positive effect, we require more combinations of the base and rotated images to achieve the highest potential. Furthermore, optimizing the learnable parameters solely with rotations is insufficient because this process only converges the current embeddings towards the prototypes and has a minimal effect on domain adaptation.
1-shot ChestX ISIC EuroSAT CropDisease Avg. $\uparrow$
CPLSR $\boldsymbol{23.00\pm 0.31}$ $\boldsymbol{33.49\pm 0.42}$ $\boldsymbol{74.54\pm 0.60}$ $\boldsymbol{83.46\pm 0.58}$ $\boldsymbol{53.62}$
CPs + Pseudo-classes $22.90\pm 0.31$ $33.39\pm 0.43$ $73.72\pm 0.61$ $82.09\pm 0.59$ $53.03$
CPs + SSTs $22.91\pm 0.32$ $33.36\pm 0.43$ $73.56\pm 0.61$ $82.11\pm 0.59$ $52.99$
DINO $22.92\pm 0.32$ $33.24\pm 0.44$ $73.59\pm 0.61$ $82.01\pm 0.59$ $52.94$
AttnScale $22.87\pm 0.32$ $33.28\pm 0.44$ $73.58\pm 0.60$ $81.93\pm 0.59$ $52.92$
Plain Prompts,length=2 $22.79\pm 0.33$ $32.66\pm 0.43$ $72.53\pm 0.62$ $81.96\pm 0.59$ $52.49$
5-shot ChestX ISIC EuroSAT CropDisease Avg. $\uparrow$
CPLSR $\boldsymbol{27.14\pm 0.33}$ $\boldsymbol{48.08\pm 0.47}$ $\boldsymbol{90.69\pm 0.31}$ $\boldsymbol{95.11\pm 0.31}$ $\boldsymbol{65.26}$
CPs + Pseudo-classes $26.69\pm 0.33$ $47.90\pm 0.46$ $89.96\pm 0.31$ $94.68\pm 0.32$ $64.81$
DINO $26.86\pm 0.33$ $47.87\pm 0.46$ $89.79\pm 0.31$ $94.70\pm 0.32$ $64.81$
AttnScale $26.85\pm 0.33$ $47.87\pm 0.47$ $89.79\pm 0.31$ $94.67\pm 0.32$ $64.80$
CPs + SSTs $26.73\pm 0.33$ $47.92\pm 0.46$ $89.86\pm 0.31$ $94.66\pm 0.32$ $64.79$
Plain Prompts,length=2 $26.77\pm 0.33$ $47.02\pm 0.47$ $88.71\pm 0.34$ $94.66\pm 0.32$ $64.29$
Table 3: Ablation studies on Mini-ImageNet
6.6 UMAP analysis
(a) Mini-ImageNet - Before
(b) Mini-ImageNet - After
(c) CropDisease - Before
(d) CropDisease - After
Figure 4: UMAP graphs for the Mini-ImageNet and CropDisease datasets (only real classes).
Fig. 4 shows the UMAP [19] for the Mini-ImageNet and CropDisease datasets, before and after training. The top graphs show that since the pseudo-classes are less likely to be chosen in areas where many base domain classes exist, they would compact those areas of the latent space. The bottom graphs reveal that the condensation of the base class clusters results in stretching the latent space and enhancing the separation of the target domain classes.
The interactions or dynamics between the embeddings of the rotated images plus the real images’ embeddings on the UMAP graphs might be complicated, as there are many classes. Besides, if we want to observe the base and target classes at the same time, the number of classes will explode. Therefore, it requires a comprehensive future research to study the other possible effects or nuances of applying CPLSR.
7 Discussion and Future Opportunities
Despite the advantages of the CPLSR, some possible drawbacks may require further research in the future. First, since the space occupied by the base classes in the latent space is being compressed, this process may decrease the network’s ability to separate them in the base domain; consequently, accuracy may decrease when the network encounters samples from similar domains to the base. Second, in the prototypical networks, the repulsive force vanishes exponentially after the embeddings of different classes are sufficiently distinct from any specific class, due to the softmax calculations. Therefore, we see in practice that the network reaches its highest accuracy on the validation set after hundreds of episodes. Applying different loss functions with margin terms to study the effect of keeping the repulsive force further is an intriguing field to explore.
Here, we observe the CP between the query and key tensors in the attention (after two projections). It has the potential to be used in other parts of unimodal or multimodal models to combine two concepts into a single concept or replace operations involving two linear projections anywhere.
8 Conclusion
Our finding shows that current methods underperform DINO in CD-FSL. This paper proposes CPLSR. It consists of the Coalescent Projection (CP)and Latent Space Reservation (LSR)to outperform DINO for the first time. The CP is the successor to the plain/soft prompts, which works as a unified projection with a single matrix, rather than having two projections. Moreover, the Latent Space Reservation (LSR) reserves the latent space and stretches it by generating novel pseudo-classes in both the input and embedding spaces of the base domain to anticipate new, unseen samples from extreme domain shift scenarios. Our rigorous numerical studies demonstrate that our approach outperforms prior arts, including DINO, by notable margins. Our future work is devoted to addressing the problem of few-shot defect classifications.

### v2

@@ 6 Conclusion
To analyze the third term of the denominator, Eq. 46, consider classes $c_{1}$ and $c_{2}$. This term is therefore becoming ${\operatorname{E}}[(\parallel\mu_{c_{1}}-\mu_{c_{2}}\parallel^{2})^{2}]={\operatorname{E}}[\parallel\mu_{c_{1}}-\mu_{c_{2}}\parallel^{4}]$. While it grows, there are two mechanisms that limit its growth. Firstly, due to the usage of the cosine distance, we know $\parallel\mu_{c}\parallel<=1$, hence $\parallel\mu_{c_{1}}-\mu_{c_{2}}\parallel\leq 2$, and $\parallel\mu_{c_{1}}-\mu_{c_{2}}\parallel^{4}<16$. Thus, it cannot be exploded. Secondly, due to the novel-novel and base-novel filters, the outliers are being filtered, which curbs the extreme growth of pairwise distances.
Overall, the numerator increases while $\operatorname{V}_{\operatorname{Tr}}(\Sigma_{y})$ and $\operatorname{V}_{\mathrm{wit}}(\Sigma_{\tau},\Sigma,\mu)$ as the dominant terms of the denominator decrease and the fraction increases.
In Eq. 49 we know from [ben2010theory] that $\hat{d}_{H\Delta H}(D_{S},D_{T})=2\operatorname{sup}_{h,h^{\prime}\in H}|\operatorname{pr}_{x\sim D_{S}}[h(x)\neq h^{\prime}(x)]-\operatorname{pr}_{x\sim D_{T}}[h(x)\neq h^{\prime}(x)]|$, which measures how easily hypotheses in the hypothesis space $H$ can separate the source from the target using their disagreement regions. The pseudo-embeddings fill previously unoccupied parts of the latent space and increase the overlap between the subsets of $H$. Therefore, the domain discriminator in $H$ sees closer occupancy patterns and would have less power, leading to the reduction in the empirical $H\Delta H$}-divergence.
Regarding the ideal joint hypothesis $\lambda=R_{B}(h^{*})+R_{T}(h^{*})$, where $h^{*}=\arg\min_{h\in H}(R_{B}(h)+R_{T}(h))$, LSR bends cosine embeddings to be closer to the prototypes, and decision regions are supported by larger margins. This process increases the probability of finding a hypothesis from $H$ that attains lower error on both domains, despite belonging to the different classes, hence the $\lambda$ will be reduced.
At last, in the final term of the Eq. 49, since LSR increases the amount and samples of the source domain, the $m$ becomes larger, which reduces the risk.
Appendix S4 Pseudo-code
The pseudo-code of the CPLSR is presented in Algorithm 1.
Appendix S5 Time Complexity Analysis
@@
The crucial parts of a ViT are the Multi-Head Self-Attention (MHSA), Multi-Layer Perceptrons (MLPs) in the transformer blocks, and the patch embedding convolution. A convolution layer is equivalent to a linear layer and is negligible in comparison to the overall calculations in $L$ blocks. The patch embedding layer generates $n=\frac{HW}{P^{2}}+1$ tokens (1 for the CLS token), where $H$, $W$, and $P$ are the image height, width, and patch size, respectively
The $Q$, $K$, VVprojections require $O(nd^{2})$, where $d=n_{h}d_{k}$. $QK^{\text{T}}$ requires $O(n_{h}n^{2}d_{h})=O(n^{2}d)$ operations. $\operatorname{Softmax}$ costs $O(n_{h}n^{2})$. Multiplication to $V$ requires additional $O(n_{h}n^{2}d_{k})=O(n^{2}d)$ operation. The output projection requires $O(nd^{2})$ operations. Finally, the MLP for output projection requires $O(L_{\text{MLP}}ndd_{\text{MLP}})$. Since the MLP is shallow in practice ($1\leq L_{\text{MLP}}\leq 2$), it requires $O(ndd_{\text{MLP}})$. In conclusion, a vanilla MHSA costs $O(nd^{2}+n^{2}d+ndd_{\text{MLP}})$.
With CP, we have $QCK^{\text{T}}$ instead of $QK^{\text{T}}$. The cost of $QC$ is $O(n_{h}nd_{k}^{2})=O(\frac{nd^{2}}{n_{h}})$, and the multiplication with $K$ requires additional $O(n_{h}n^{2}d_{k})=O(n^{2}d)$. Thus, the time complexity for $QCK^{\text{T}}$ is $O(\frac{nd^{2}}{n_{h}}+n^{2}d)$. Finally, the $\operatorname{Softmax}(\frac{QCK^{\text{T}}}{\sqrt{d_{k}}})V$ requires $O(\frac{nd^{2}}{n_{h}})+n^{2}d)+O(n^{2}d)=O(nd^{2}+n^{2}d)$. Therefore, the overall complexity of an MHSA block is $O(nd^{2}+n^{2}d+ndd_{\text{MLP}})$, which is the same as the time-complexity of the vanilla attention module.
@@
Overall, the total time-complexity of a ViT with $L$ transformer blocks is $O\left(L(dn^{2}+nd^{2}+ndd_{\text{MLP}})\right)$.
Appendix S6 CP vs. AttnScale
Mahalanobis distance metric: While CP is unconstrained by default, it has the ability to be constrained to learn with the Mahalanobis distance metric, while AttnScale does not have such an ability. Mahalanobis distance is shown that is able to improve the performance in FSL and prototype-based methods, such as Simple CNAPS [bateni2020improved] and ROBUSTA [Paeedeh2024FewShotCI]. Moreover, Eq. (12) in the main paper utilizes a squared of the Mahalanobis distance term: $(\mu_{\tilde{k}}-\mu_{k})^{\text{T}}\Sigma_{\tilde{k}}^{\dagger}(\mu_{\tilde{k}}-\mu_{k}))$ to assign base-novel scores, which our learnable parameters (CP) should have a potential to learn it. Finally, using the dot-product attention results in a hyper-spherical neighborhood around each query [nielsen2024elliptical]. If the mapping varies along different directions, dealing with all coordinates as equally important would be suboptimal. Therefore, learning a Mahalanobis distance is crucial. Mahalanobis distance can be written as [kulis2013metric, weinberger2009distance]:
$d_{\mathrm{Mah}}(x_{i},x_{j})=\sqrt{(x_{i}-x_{j})^{\text{T}}\Sigma^{-1}(x_{i}-x_{j})},$ (50)
where $\Sigma$ is the positive semi-definite covariance matrix.
To utilize the Mahalanobis distance in the neighborhood of the query as:
$d_{\mathrm{Mah}}(q,k):=\sqrt{(q-k)^{\text{T}}C(q-k)},$ (51)
we can define our learnable CP as $C=PP^{\text{T}}$ with Cholesky decomposition. Moreover, to maintain the symmetry, we can use the symmetry regularizer
$\lambda\parallel C-C^{\text{T}}\parallel_{F}^{2}$ (52)
during the training, where $\lambda$ is a small scalar as a constraint to guide the CP to learn the Mahalanobis distances.
Moreover, we can deal with Mahalanobis distance as an Euclidean distance after applying a linear projection [kulis2013metric]. If $G=P^{\text{T}}$:
$d_{\mathrm{Mah}}(q,k)$ $=\sqrt{(q-k)^{\text{T}}C(q-k)}$ (53)
$=\sqrt{(q-k)^{\text{T}}G^{\text{T}}G(q-k)}$ (54)
$=\sqrt{\Bigl(G(q-k)\Bigr)^{\text{T}}\Bigl(G(q-k)\Bigr)}$ (55)
$=\parallel(G(q-k)\parallel_{2}$ (56)
$=\parallel(Gq-Gk)\parallel_{2}$ (57)
$=d_{\mathrm{Euc}}(Gq,Gk).$ (58)
In AttnScale, since it reweights the token-to-token pairs rather than feature dimensions, it cannot parameterize a Mahalanobis metric in the latent space; hence, it cannot stretch or shrink in the latent space.
Appendix S7 Ablation Study
In the main paper, we showed the one soft prompt length in the ablation study subsection due to the page limit. Here, we present additional experiments with varying prompt lengths in Tab. 4, which could not be covered in the main paper.
As the results show, plain prompts in none of the cases could outperform DINO. Moreover, increasing the prompt length has a detrimental effect on the challenging task of CD-FSL, as the number of interactions between tokens increases from the first to the final layers.